# A Comprehensive Review on Harnessing Large Language Models to Overcome Recommender System Challenges

### v1

@@ 2. Evolution of Architectures in Recommender Systems
The architectural landscape of recommender systems has undergone a remarkable transformation over the past two decades. Beginning with simple heuristic methods and collaborative filtering techniques, the field has progressively embraced more sophisticated paradigms powered by deep learning, graph neural networks, and, more recently, LLMs. This evolution reflects the growing demand for personalization, scalability, and contextual awareness in real-world applications. This section presents a chronological overview of key architectural advancements in recommender systems, highlighting the motivations, technical foundations, and practical implications of each generation. By tracing this progression, we aim to shed light on how the field has adapted to increasingly complex data modalities, user behaviors, and computational constraints.
2.1. Traditional Approaches
Heuristic-Based: Heuristic-based recommender systems represent some of the earliest strategies for generating item suggestions. These approaches rely on deterministic rules and statistical aggregations derived from historical data, rather than modeling latent user-item interactions. They are often used in large-scale systems for bootstrapping recommendations or serving as baseline models (Schafer et al., 2001; Bobadilla et al., 2018). A common heuristic is popularity-based recommendation, which ranks items based on global interaction frequency. For instance, the recommendation score $s_{i}$ for an item $i$ may be computed as:
$s_{i}=\text{count}(i)=\sum_{u\in\mathcal{U}}\mathbb{I}(i\in I_{u})$
where $\mathbb{I}(\cdot)$ is the indicator function, $I_{u}$ is the set of items interacted with by user $u$, and $\mathcal{U}$ is the set of all users. This formulation favors frequently interacted items, such as most-viewed videos or best-selling products.
Another widely used rule is co-occurrence-based recommendation, which suggests items based on observed pairwise co-interactions. For example, an associative rule like “Users who bought or viewed item $i$ also bought or viewed item $j$” is operationalized using:
$\text{sim}(i,j)=\frac{\text{count}(i,j)}{\sqrt{\text{count}(i)\cdot\text{count}(j)}}$
where $\text{count}(i,j)$ is the number of users who interacted with both items $i$ and $j$, serving as a symmetric co-occurrence similarity metric (Sarwar et al., 2001). Despite their simplicity and scalability, heuristic methods are inherently non-personalized—they assume user homogeneity and do not adapt to individual preferences. As such, they suffer from popularity bias and lack context-awareness, making them less suitable for tasks requiring fine-grained personalization (Adomavicius and Tuzhilin, 2005). Nonetheless, their low computational cost and interpretability make them useful in cold-start scenarios where user history is sparse, in bootstrapping candidate sets in multi-stage recommender pipelines, and in serving fallback recommendations when more advanced model outputs are unreliable. In many production systems, heuristic scores are combined with learned models in later ranking stages or incorporated into ensemble frameworks for robustness.
Figure 1. (i) Content-Based Recommendation: Recommends movies based on content features such as genre and cast similarity with previously watched movies. (ii) User-Based Collaborative Filtering: Recommends movies by identifying users with similar tastes and suggesting items liked by similar users. (Mallick, 2023)
Collaborative Filtering: Collaborative Filtering (CF)  (Zhang et al., 2014) is a core methodology in recommender systems that predicts a user’s interest in an item by leveraging patterns of user-item interactions. The central assumption behind CF is the homophily principle: users who have shown similar interaction patterns in the past are likely to share future preferences. Unlike content-based methods, CF requires no side information or item/user metadata—it exclusively relies on interaction history, such as ratings, clicks, or purchases.
Formally, given a sparse interaction matrix $\mathbf{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{I}|}$ where $\mathcal{U}$ is the set of users and $\mathcal{I}$ is the set of items, the goal is to estimate the unobserved ratings or affinities $\hat{r}_{u,i}$ for a user-item pair $(u,i)$.
There are two major paradigms of CF. User-based CF computes the similarity between the target user $u$ and other users $v\in\mathcal{U}$, and then aggregates the preferences of the $k$ most similar users $\mathcal{N}_{k}(u)$ to predict a score using the formula:
$\hat{r}_{u,i}=\frac{\sum_{v\in\mathcal{N}_{k}(u)}\text{sim}(u,v)\cdot r_{v,i}}{\sum_{v\in\mathcal{N}_{k}(u)}|\text{sim}(u,v)|}$
where $\text{sim}(u,v)$ may be computed using cosine similarity, Pearson correlation, or Jaccard similarity over the vectors $\mathbf{r}_{u}$ and $\mathbf{r}_{v}$.
In contrast, item-based CF finds items similar to those the user has interacted with, using item-item similarities. The estimated score is:
$\hat{r}_{u,i}=\frac{\sum_{j\in\mathcal{I}_{u}}\text{sim}(i,j)\cdot r_{u,j}}{\sum_{j\in\mathcal{I}_{u}}|\text{sim}(i,j)|}$
where $\mathcal{I}_{u}$ is the set of items interacted with by user $u$, and similarity can again be cosine or adjusted cosine similarity.
Common similarity functions include cosine similarity:
$\text{sim}_{\text{cosine}}(i,j)=\frac{\mathbf{r}_{*,i}\cdot\mathbf{r}_{*,j}}{\|\mathbf{r}_{*,i}\|\cdot\|\mathbf{r}_{*,j}\|}$
and Pearson correlation:
$\text{sim}_{\text{Pearson}}(i,j)=\frac{\sum_{u}(r_{u,i}-\bar{r}_{i})(r_{u,j}-\bar{r}_{j})}{\sqrt{\sum_{u}(r_{u,i}-\bar{r}_{i})^{2}}\sqrt{\sum_{u}(r_{u,j}-\bar{r}_{j})^{2}}}$
Despite its simplicity and interpretability, CF faces several challenges. Chief among them are data sparsity and the cold-start problem—both user and item-level—which hinder similarity computation. Real-time similarity computation can also become computationally expensive as the number of users or items grows. To mitigate these limitations, CF is often combined with content-based models in hybrid architectures.
Furthermore, recent advances have proposed learning-based CF methods such as matrix factorization and neural embeddings. For instance, LightGCN (He et al., 2020) simplifies graph-based collaborative filtering by removing feature transformations and non-linearities, using only weighted summation of neighborhood embeddings:
$\mathbf{e}_{u}^{(k)}=\sum_{i\in\mathcal{N}(u)}\frac{1}{\sqrt{|\mathcal{N}(u)||\mathcal{N}(i)|}}\cdot\mathbf{e}_{i}^{(k-1)}$
This iterative propagation enhances higher-order connectivity, improving recommendation accuracy while maintaining linear complexity. Collaborative Filtering remains a fundamental building block in many industrial pipelines and serves as a baseline for benchmarking modern deep learning recommender systems (Li et al., 2023d; Rendle et al., 2020).
Matrix Factorization: Matrix Factorization (MF)  (Zhang, 2022) is a class of latent factor models that projects users and items into a shared low-dimensional vector space to model user-item interactions. Given a user–item rating matrix $R\in\mathbb{R}^{m\times n}$, where $m$ is the number of users and $n$ is the number of items, the goal is to approximate $R$ by the product of two low-rank matrices:
$R\approx\hat{R}=UV^{\top},$
where $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$ are the user and item latent factor matrices, and $k\ll\min(m,n)$ is the number of latent dimensions.
The optimization objective is typically to minimize the squared error over observed entries:
$\min_{U,V}\sum_{(u,i)\in\mathcal{K}}(r_{u,i}-\mathbf{u}_{u}^{\top}\mathbf{v}_{i})^{2}+\lambda\left(\|\mathbf{u}_{u}\|^{2}+\|\mathbf{v}_{i}\|^{2}\right),$
where $\mathcal{K}$ is the set of observed ratings, $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the latent vectors for user $u$ and item $i$, respectively, and $\lambda$ is a regularization hyperparameter to prevent overfitting.
Several variations of matrix factorization have been proposed to enhance recommendation accuracy and scalability. Singular Value Decomposition (SVD)  (Klema and Laub, 1980) applies truncated decomposition to the user-item rating matrix, projecting it into lower-dimensional latent factors; however, it often requires prior imputation of missing entries, which can introduce bias. Alternating Least Squares (ALS) offers a more scalable alternative by optimizing the matrix factorization objective through iterative updates: fixing one of the user or item matrices while solving a regularized least squares problem for the other (Zhou et al., 2008). This alternating procedure makes ALS particularly well-suited for large-scale and distributed environments, such as those in industrial settings. An extension to these approaches, SVD++, incorporates implicit feedback—such as clicks or views—alongside explicit ratings to refine user representations. The predicted rating in SVD++ is given by:
$\hat{r}_{u,i}=\mu+b_{u}+b_{i}+\left(\mathbf{u}_{u}+\sum_{j\in\mathcal{I}_{u}}\frac{y_{j}}{\sqrt{|\mathcal{I}_{u}|}}\right)^{\top}\mathbf{v}_{i},$
where $\mu$ is the global average rating, $b_{u}$ and $b_{i}$ are user and item-specific biases, $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the latent factors for user $u$ and item $i$, respectively, and $y_{j}$ represents the embedding of item $j$ contributing to user $u$’s implicit feedback profile.
Matrix factorization has become the backbone of many collaborative filtering systems due to its balance between expressiveness and scalability. Its popularity surged during the Netflix Prize competition, and it continues to serve as a robust baseline for collaborative filtering tasks (Koren et al., 2009).
2.2. Content-Based and Hybrid Systems
Content-Based Filtering: Content-based filtering (CBF)  (Wang et al., 2024d) recommends items to a user by analyzing the similarity between the items’ attributes and the user’s historical preferences (Lops et al., 2011). Unlike collaborative methods, CBF operates independently of other users and relies solely on item metadata—such as genre, tags, textual descriptions, or visual features.
A user profile $\mathbf{u}$ is typically constructed by aggregating the feature vectors of previously liked or interacted items. For a given item $i$ with a feature vector $\mathbf{x}_{i}$, the predicted preference score is computed using a similarity function:
$\hat{r}_{u,i}=\text{sim}(\mathbf{u},\mathbf{x}_{i})$
Common choices for sim include cosine similarity:
$\text{sim}_{\text{cosine}}(\mathbf{u},\mathbf{x}_{i})=\frac{\mathbf{u}^{\top}\mathbf{x}_{i}}{\|\mathbf{u}\|\cdot\|\mathbf{x}_{i}\|}$
In text-based domains, TF-IDF representations or pre-trained language model embeddings are often used to define $\mathbf{x}_{i}$, while in visual domains, image embeddings from CNNs or vision transformers are leveraged (Mouhiha et al., 2024).
While CBF models are robust to cold-start users (as long as item features are known), they struggle with over-specialization and limited discovery of novel content.
Hybrid Models: Hybrid recommender systems combine content-based and collaborative filtering paradigms to overcome the limitations of either approach in isolation (Burke, 2002). The fusion can occur at various levels—model, feature, or score—and often improves robustness, especially in sparse or cold-start scenarios.
LightFM (Kula, 2015) incorporates both collaborative signals (user-item interactions) and content features (metadata) into a unified latent factor model. Its objective is to learn user and item embeddings from both sources:
$\hat{r}_{u,i}=\mathbf{u}_{u}^{\top}\mathbf{v}_{i}+\mathbf{u}_{u}^{\top}\mathbf{f}_{i}+\mathbf{f}_{u}^{\top}\mathbf{v}_{i},$
where $\mathbf{f}_{u}$ and $\mathbf{f}_{i}$ are feature representations of user and item metadata respectively.
DeepFM (Guo et al., 2017) further extends this by combining a Factorization Machine (FM) layer for low-order feature interactions with a deep neural network for high-order interactions. while the deep component captures non-linear patterns across raw features via multi-layer perceptrons. These hybrid architectures provide stronger generalization across heterogeneous data sources and are widely used in large-scale recommendation engines.
2.3. Deep Learning Architectures
Neural Collaborative Filtering (NCF): Neural Collaborative Filtering (NCF)  (He et al., 2017a) is a deep learning-based extension of matrix factorization that captures complex, nonlinear user–item interaction patterns. Instead of modeling the dot product between latent user and item vectors, NCF uses a Multi-Layer Perceptron (MLP) to learn a nonlinear interaction function:
$\hat{r}_{u,i}=\phi(\mathbf{u}_{u},\mathbf{v}_{i})=\text{MLP}([\mathbf{u}_{u};\mathbf{v}_{i}]),$
where $\mathbf{u}_{u}$ and $\mathbf{v}_{i}$ are the embedding vectors for user $u$ and item $i$, respectively, and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation. The architecture enables the model to approximate arbitrary functions between user and item representations. The NeuMF model (He et al., 2017b) combines Generalized Matrix Factorization (GMF) and MLP-based modeling via a fusion layer to benefit from both linear and nonlinear signals.
Figure 2. Neural Collaborative Filtering (NCF) architecture: One-hot user and item inputs are embedded into latent vectors, passed through neural layers, and produce a predicted score $\hat{y}_{ui}$ trained against the target $y_{ui}$ (He et al., 2017a)
Wide & Deep Models: The Wide & Deep architecture (Cheng et al., 2016) aims to capture both low- and high-order feature interactions by combining a linear model (”wide”) and a deep neural network (”deep”) within a joint framework:
$\hat{y}=\sigma(\mathbf{w}^{\top}\mathbf{x}+\mathbf{a}^{\top}f_{\text{deep}}(\mathbf{x})),$
where $\mathbf{x}$ is the feature vector, $\mathbf{w}$ are the wide (memorization) weights, and $f_{\text{deep}}(\mathbf{x})$ represents the output of the deep network. This hybrid design enables the model to both remember frequent co-occurrences (e.g., user clicked item A and B) and generalize to unseen feature combinations. It was deployed in production at Google Play for app recommendations, demonstrating substantial improvements in personalization and click-through rates.
2.4. Sequence-Aware Models
As user behavior in many applications is inherently sequential (e.g., streaming, shopping, or reading), capturing temporal dependencies becomes crucial. Sequence-aware recommender systems model the order and dynamics of user-item interactions over time, enabling next-item prediction based on behavioral context.
RNN-Based Models: Recurrent Neural Networks (RNNs)  (Schmidt, 2019), particularly Gated Recurrent Units (GRUs) and Long Short-Term Memory (LSTM) networks, have been used to model user behavior sequences. GRU4Rec (Hidasi et al., 2016) was one of the earliest models to apply RNNs to session-based recommendation, treating the problem as a sequence prediction task.
Let $x_{t}$ be the embedding of the item interacted with at time $t$. The GRU updates its hidden state $h_{t}$ using:
$h_{t}=\text{GRU}(x_{t},h_{t-1}),$
and predicts the next item via:
$\hat{y}_{t+1}=\text{softmax}(Wh_{t}+b),$
where $W$ and $b$ are learnable parameters. GRU4Rec enables learning from both short-term (within-session) and long-term user preferences and has shown superior performance in session-based scenarios compared to traditional CF methods.
Transformer-Based Models: Transformer-based  (Vaswani et al., 2023) models replace the recurrence in RNNs with self-attention mechanisms, allowing better modeling of long-range dependencies and parallelization. SASRec (Kang and McAuley, 2018) applies a unidirectional self-attention encoder to model sequential interactions:
$\mathbf{H}=\text{SelfAttention}(\mathbf{X}+\mathbf{P}),$
where $\mathbf{X}$ is the sequence of item embeddings and $\mathbf{P}$ is the positional encoding. The model predicts the next item using the last hidden state:
$\hat{y}_{t+1}=\text{softmax}(Wh_{t}).$
BERT4Rec (Sun et al., 2019) introduces bidirectional attention with a masked item prediction objective, inspired by the BERT architecture in NLP. This enables the model to learn contextual dependencies from both past and future items in a sequence.
TiSASRec (Li et al., 2020) extends SASRec by incorporating both absolute and relative time intervals into attention weights, enhancing temporal modeling in user behavior:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}+T_{\text{abs}}+T_{\text{rel}}}{\sqrt{d_{k}}}\right)V,$
where $T_{\text{abs}}$ and $T_{\text{rel}}$ encode absolute and relative temporal distances, respectively.
Overall, transformer-based models have surpassed RNNs in many sequence-aware recommendation tasks due to their ability to model long-term dependencies and leverage parallel computation.
2.5. Graph-Based Recommenders
Graph-based recommender systems  (Wang et al., 2021a) represent user–item interactions as a bipartite graph $\mathcal{G}=(\mathcal{U}\cup\mathcal{I},\mathcal{E})$, where $\mathcal{U}$ is the set of users, $\mathcal{I}$ is the set of items, and edges $\mathcal{E}$ correspond to observed interactions (e.g., clicks, ratings, purchases). This graph structure enables the modeling of high-order connectivity and collaborative signals beyond immediate neighbors, capturing richer semantics than traditional matrix factorization.
Figure 3. Illustration of a heterogeneous graph-based recommender system integrating three components: the Social Relation Graph (user-user links), the User-Item Interaction Graph (explicit and implicit feedback such as clicks and ratings), and the Domain Knowledge Graph (item attributes like OS and color). This structure facilitates richer context-aware and socially-informed recommendations (Kipf and Welling, 2017)
Graph Convolutional Networks (GCNs) have been widely adopted for such modeling  (Kipf and Welling, 2017). A GCN propagates information between nodes through message passing, updating the representation of a node by aggregating features from its neighbors. In a simplified GCN layer:
$\mathbf{h}_{v}^{(k)}=\sigma\left(\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{W}^{(k)}\mathbf{h}_{u}^{(k-1)}\right),$
where $\mathbf{h}_{v}^{(k)}$ is the embedding of node $v$ at the $k^{\text{th}}$ layer, $\mathbf{W}^{(k)}$ is a learnable weight matrix, and $\mathcal{N}(v)$ is the set of neighbors of $v$.
NGCF (Neural Graph Collaborative Filtering) (Ying et al., 2018) builds on this idea by combining collaborative filtering with graph convolution. It explicitly models user–item interactions via non-linear message propagation. At each layer, it computes:
$\mathbf{e}_{v}^{(k)}=\sigma\left(\mathbf{W}_{1}^{(k)}\mathbf{e}_{v}^{(k-1)}+\mathbf{W}_{2}^{(k)}\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{e}_{u}^{(k-1)}\right),$
where $\mathbf{e}_{v}^{(k)}$ is the embedding of user or item $v$ at layer $k$.
LightGCN (He et al., 2020) simplifies NGCF by removing the nonlinear activation and transformation weights, arguing that they may over-complicate the propagation. LightGCN computes embeddings as:
$\mathbf{e}_{v}=\sum_{k=0}^{K}\alpha_{k}\mathbf{e}_{v}^{(k)},\quad\text{where}\quad\mathbf{e}_{v}^{(k)}=\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{|\mathcal{N}(v)||\mathcal{N}(u)|}}\mathbf{e}_{u}^{(k-1)}.$
PinSage (Ying et al., 2018) extends this concept for large-scale web-scale recommendation systems (e.g., Pinterest). It combines random walks with graph convolutions and incorporates node features using neighborhood sampling and importance-based aggregation.
Graph-based models are especially effective at capturing high-order collaborative signals and performing inductive reasoning on unseen nodes. They have become foundational in modern recommenders deployed in large-scale platforms due to their flexibility and scalability.
2.6. Large-Scale Retrieval Models
Large-scale retrieval models are designed to efficiently handle recommendation tasks at industrial scale, where the corpus of items may reach millions or billions. These models prioritize fast candidate generation and scalability while maintaining reasonable recommendation quality.
Figure 4. Hybrid ranking architecture combining two-tower and pairwise models. User-specific item representations and shallow embeddings are processed through fusion scoring to generate the final ranked list.  (Yuan et al., 2024).
Two-Tower Models: Two-tower architectures, also known as dual-encoder models  (Yuan et al., 2024), are among the most commonly deployed retrieval methods in production systems. These models learn independent embedding functions for users and items:
$\mathbf{e}_{u}=f_{\theta}(u),\quad\mathbf{e}_{i}=g_{\phi}(i)$
where $f_{\theta}$ and $g_{\phi}$ are neural networks parameterized by $\theta$ and $\phi$, respectively. The final relevance score between user $u$ and item $i$ is computed as the inner product or cosine similarity:
$s(u,i)=\langle\mathbf{e}_{u},\mathbf{e}_{i}\rangle$
This separation allows precomputing item embeddings and indexing them via Approximate Nearest Neighbor (ANN) techniques, enabling sublinear retrieval times.
The YouTube DNN model (Covington et al., 2016) is an influential example of a two-tower retrieval system. It uses multi-layer perceptrons (MLPs) for both user and item towers, trained with sampled softmax to optimize for the target watch history. Another widely used architecture is Facebook’s Deep Learning Recommendation Model (DLRM) (Naumov et al., 2019), which combines dense and sparse features from user-item pairs and leverages embedding lookups and interaction layers to compute the final relevance score.
These models form the first stage in multi-stage ranking pipelines. After candidate generation via ANN search over user-item embeddings, a more expressive model (e.g., gradient boosting, transformers) typically performs re-ranking over the retrieved shortlist.
Due to their efficiency and deployment advantages, two-tower retrieval models are considered foundational in modern industrial recommender systems.

### v2

@@ 1. Introduction
Figure 1. Taxonomy of recommender system challenges and corresponding LLM-based solutions. The framework categorizes challenges into four major areas—data-centric, modeling and algorithmic, evaluation, and privacy/security—each with sub-problems and LLM-driven strategies for addressing them.’=
@@ 2. Evolution of Architectures in Recommender Systems
Recommender systems have progressed from heuristic rules to collaborative filtering, latent factorization, neural and graph-based models, and finally industrial-scale retrieval frameworks. Each paradigm enhanced expressiveness, scalability, and the ability to capture complex dependencies.
2.1. From Heuristics to Latent Factor Models
Early recommenders used deterministic heuristics such as popularity- and co-occurrence-based rules (Schafer et al., 2001; Bobadilla et al., 2018; Sarwar et al., 2001), which scale well but lack personalization. Collaborative Filtering (CF) (Zhang et al., 2014) introduced personalization by exploiting similarity in interaction patterns between users or items. While simple and interpretable, CF suffers from sparsity and cold-start issues.
Matrix Factorization (MF) (Klema and Laub, 1980; Koren et al., 2009) addressed scalability by embedding users and items into a shared low-dimensional space, with predictions computed via inner products of latent vectors. Variants such as ALS (Zhou et al., 2008) and SVD++ (Koren et al., 2009) incorporated implicit feedback and became widely adopted in industrial systems due to their balance of accuracy and efficiency.
2.2. Hybrid and Neural Extensions
Content-Based Filtering (CBF) (Lops et al., 2011) enriched recommendations with item attributes such as textual, categorical, or visual embeddings, making it more robust to cold-start users. However, CBF tends to overspecialize. Hybrid models (Burke, 2002) combined collaborative and content signals, exemplified by LightFM (Kula, 2015) and DeepFM (Guo et al., 2017), which integrate metadata with learned embeddings.
Deep learning extended this line with Neural Collaborative Filtering (NCF) (He et al., 2017), which replaces linear interaction functions with nonlinear neural networks, and Wide & Deep (Cheng et al., 2016), which jointly optimizes memorization and generalization. These architectures established the basis for modern industrial recommenders.
2.3. Sequential and Graph-Based Models
User behavior often exhibits temporal and contextual dynamics. Sequence-aware recommenders such as GRU4Rec (Hidasi et al., 2016) applied recurrent neural networks to session data, while transformer-based approaches (e.g., SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), TiSASRec (Li et al., 2020)) leveraged self-attention to capture long-range dependencies and temporal intervals. These models outperform traditional CF in sequential domains such as streaming and e-commerce.
In parallel, graph-based recommenders (Wang et al., 2021a) represented interactions as bipartite graphs. Neural Graph Collaborative Filtering (NGCF) (Ying et al., 2018) propagated embeddings through nonlinear message passing, while LightGCN (He et al., 2020) simplified this to linear aggregation, improving both accuracy and scalability. PinSage (Ying et al., 2018) scaled graph-based recommendation to billions of nodes using random walks and sampling, enabling deployment in web-scale platforms like Pinterest.
2.4. Industrial-Scale Retrieval Frameworks
With item catalogs often exceeding billions, retrieval efficiency became a dominant concern. Two-tower architectures (Yuan et al., 2024) learn independent embeddings for users and items, with relevance computed as the inner product between vectors. Pre-computed item embeddings can be indexed using Approximate Nearest Neighbor (ANN) search, enabling sublinear retrieval times.
The YouTube DNN (Covington et al., 2016) demonstrated the effectiveness of deep two-tower retrieval with sampled softmax, while Facebook’s DLRM (Naumov et al., 2019) integrated sparse and dense features with cross-feature interaction layers. These retrieval modules form the candidate generation stage in multi-stage pipelines, with re-ranking typically handled by more expressive models such as gradient boosting or transformers.